AN EXAMPLE OF A DOMAIN WITH NON-COMPACT AUTOMORPHISM GROUP

SIQI FU ALEXANDER V. ISAEV STEVEN G. KRANTZ

Abstract. We give an example of a bounded, pseudoconvex, circular domain in ℂ^{3} with smooth, real-analytic boundary and non-compact automorphism group, which is not biholomorphically equivalent to any Reinhardt domain.